# Leveraging non-unital noise for gate-based quantum reservoir computing

### v1

@@ Appendix A Contractivity of noise channels
$\mathcal{D}_{\text{bf}}(\rho)=\sqrt{1-\gamma}\rho+\sqrt{\gamma}X\rho X.$
@@
$\mathcal{D}_{\text{bpf}}(\rho)$ $=\sqrt{1-\gamma}\rho+\sqrt{\gamma}Y\rho Y.$
$\mathcal{D}_{\text{pf}}(\rho)$ $=\sqrt{1-\gamma}\rho+\sqrt{\gamma}Z\rho Z.$
@@
$K_{0}=\sqrt{1-\gamma}\,\mathbb{I}\mbox{\quad and\quad}K_{1}=\sqrt{\gamma}\,X$
@@
$\norm{K_{0}(\rho-\sigma)K_{0}^{\dagger}}_{1}\leq 2\left(1-\gamma\right)|\rho_{00}-\sigma_{00}|$
@@
$\norm{K_{1}(\rho-\sigma)K_{1}^{\dagger}}_{1}=0.$
@@
The quantum operation that describes energy dissipation is the amplitude damping. Recalling the Kraus decomposition of the amplitude damping channel, we have
$\mathcal{D}_{\text{ad}}(\rho)=K_{0}\rho K_{0}^{\dagger}+K_{1}\rho K_{1}^{\dagger}$
@@
$\norm{K_{0}(\rho-\sigma)K_{0}^{\dagger}}_{1}=|\rho_{00}-\sigma_{00}|\left(1+\sqrt{1-\gamma}\right)\leq$
$\leq r\norm{\rho-\sigma}=2|\rho_{00}-\sigma_{00}|$
@@
$\norm{K_{1}(\rho-\sigma)K_{1}^{\dagger}}_{1}\leq r\norm{\rho-\sigma}_{1}$
@@
$\norm{\mathcal{D}_{\text{ad}}(\rho)-\mathcal{D}_{\text{ad}}(\sigma)}_{1}\leq\norm{K_{0}(\rho-\sigma)K_{0}^{\dagger}}_{1}+$
$+\norm{K_{1}(\rho-\sigma)K_{1}^{\dagger}}_{1}\leq r\norm{\rho-\sigma}_{1}$
@@
$\mathcal{D}_{\text{de}}(\rho)=\frac{p}{2}\mathbb{I}+(1-p)\rho.$
In the operator-sum representation we can write
$\left(1-\frac{3p}{4}\right)\rho+\frac{p}{4}\left(X\rho X+Y\rho Y+Z\rho Z\right)$
@@
$\left(1-p\right)\rho+\frac{p}{3}\left(X\rho X+Y\rho Y+Z\rho Z\right)$
@@
$\mathcal{D}_{pd}(\rho)=\widetilde{K}_{0}\rho\widetilde{K}_{0}+\widetilde{K}_{1}\rho\widetilde{K}_{1}$
@@
$\begin{split}\widetilde{K}_{0}&=\matrixquantity(1\hfil&0\\ 0&\sqrt{1-\lambda})\\ \widetilde{K}_{1}&=\matrixquantity(0\hfil&0\\ 0&\sqrt{\lambda})\\ \end{split}$
@@
$\begin{split}K_{0}&=\sqrt{1-\gamma}\,\mathbb{I}\\ K_{1}&=\sqrt{\gamma}\,X\\ \end{split}.$
@@
Thermal relaxation is equivalent to phase-amplitude damping. The problem is that the Qiskit implementation of the thermal relaxation channel is different from the standard one, hence a short comment is needed. Thermal relaxation, in the Qiskit version, is modeled by the following Kraus decomposition in Qiskit [39]:
$\begin{split}K_{0}&=\sqrt{1-p_{z}-p_{\rm r0}-p_{\rm r1}}\,\mathbb{I}\\ K_{1}&=\sqrt{p_{z}}\,Z\\ K_{2}&=\sqrt{p_{\rm r0}}\,\matrixquantity(1\hfil&0\\ 0&0)\\ K_{3}&=\sqrt{p_{\rm r1}}\,\matrixquantity(0\hfil&0\\ 0&1)\end{split}$

### v2

@@ III Contractivity of noise channels
$\mathcal{D}_{\text{bf}}(\rho)=\sqrt{1-\gamma}\rho+\sqrt{\gamma}X\rho X.$ (23)
@@
$\mathcal{D}_{\text{bpf}}(\rho)$ $=\sqrt{1-\gamma}\rho+\sqrt{\gamma}Y\rho Y.$ (24)
$\mathcal{D}_{\text{pf}}(\rho)$ $=\sqrt{1-\gamma}\rho+\sqrt{\gamma}Z\rho Z.$ (25)
@@
$K_{0}=\sqrt{1-\gamma}\,\mathbb{I}\mbox{\quad and\quad}K_{1}=\sqrt{\gamma}\,X$ (26)
@@
$\norm{K_{0}(\rho-\sigma)K_{0}^{\dagger}}_{1}\leq 2\left(1-\gamma\right)|\rho_{00}-\sigma_{00}|$ (27)
@@
$\norm{K_{1}(\rho-\sigma)K_{1}^{\dagger}}_{1}=0.$ (28)
@@
The quantum operation that describes energy dissipation is the amplitude damping quantum channel. Recalling the Kraus decomposition of the amplitude damping channel, we have
$\mathcal{D}_{\text{ad}}(\rho)=K_{0}\rho K_{0}^{\dagger}+K_{1}\rho K_{1}^{\dagger}$ (29)
@@
$\norm{K_{0}(\rho-\sigma)K_{0}^{\dagger}}_{1}=|\rho_{00}-\sigma_{00}|\left(1+\sqrt{1-\gamma}\right)\leq$ (30)
$\leq r\norm{\rho-\sigma}=2|\rho_{00}-\sigma_{00}|$ (31)
@@
$\norm{K_{1}(\rho-\sigma)K_{1}^{\dagger}}_{1}\leq r\norm{\rho-\sigma}_{1}$ (32)
@@
$\norm{\mathcal{D}_{\text{ad}}(\rho)-\mathcal{D}_{\text{ad}}(\sigma)}_{1}\leq\norm{K_{0}(\rho-\sigma)K_{0}^{\dagger}}_{1}+$ (33)
$+\norm{K_{1}(\rho-\sigma)K_{1}^{\dagger}}_{1}\leq r\norm{\rho-\sigma}_{1}$ (34)
@@
$\mathcal{D}_{\text{de}}(\rho)=\frac{p}{2}\mathbb{I}+(1-p)\rho.$ (35)
In the operator-sum representation, we can write
$\left(1-\frac{3p}{4}\right)\rho+\frac{p}{4}\left(X\rho X+Y\rho Y+Z\rho Z\right)$ (36)
@@
$\left(1-p\right)\rho+\frac{p}{3}\left(X\rho X+Y\rho Y+Z\rho Z\right)$ (37)
@@
$\mathcal{D}_{pd}(\rho)=\widetilde{K}_{0}\rho\widetilde{K}_{0}+\widetilde{K}_{1}\rho\widetilde{K}_{1}$ (38)
@@
$\begin{split}\widetilde{K}_{0}&=\matrixquantity(1\hfil&0\\ 0&\sqrt{1-\lambda})\\ \widetilde{K}_{1}&=\matrixquantity(0\hfil&0\\ 0&\sqrt{\lambda})\\ \end{split}$ (39)
@@
$\begin{split}K_{0}&=\sqrt{1-\gamma}\,\mathbb{I}\\ K_{1}&=\sqrt{\gamma}\,X\\ \end{split}.$ (40)
@@
Thermal relaxation is equivalent to phase-amplitude damping. The problem is that the Qiskit implementation of the thermal relaxation channel differs from the standard one, hence a short comment is needed. Thermal relaxation, in the Qiskit version, is modeled by the following Kraus decomposition in Qiskit [45]:
$\begin{split}K_{0}&=\sqrt{1-p_{z}-p_{\rm r0}-p_{\rm r1}}\,\mathbb{I}\\ K_{1}&=\sqrt{p_{z}}\,Z\\ K_{2}&=\sqrt{p_{\rm r0}}\,\matrixquantity(1\hfil&0\\ 0&0)\\ K_{3}&=\sqrt{p_{\rm r1}}\,\matrixquantity(0\hfil&0\\ 0&1)\end{split}$ (41)
@@
IV Universal approximation property of our gate-based echo state network
A family of computational architectures is universal for a certain set of functionals if, for any of these mappings, there exists an instance within the family that approximates it with arbitrary precision [38]. In particular, the universality of a family of reservoir computers for fading memory mappings relies on some general sufficient conditions [10] – generalized by some of the Authors to comprehensively include quantum reservoir computing regardless of any specific implementation [40]. Here, we prove that the gate-based quantum echo state network is universal under the action of nonunital and strictly contractive noise models, thus providing theoretical support to our findings. We begin by recalling the sufficient conditions that ensure the universality, summarized in the following theorem [40].
Theorem IV.1.
Let $\mathcal{R}$ be a family of real echo reservoir computers and let =$\mathcal{I}=\left\{u\colon\mathbb{Z}\rightarrow[0,1]\right\}$ be the set of input sequences. Assume that $\left(\mathcal{I},d_{\omega}\right)$ is a compact metric space. If the set of functionals $\mathcal{H}_{\mathcal{R}}$ associated with $\mathcal{R}:$
  • has fading memory
  • separates points in $\mathcal{I}$, namely for any pair $u,v\in\mathcal{I}$ with $u\not=v$ there exists a functional $H\in\mathcal{H}_{\mathcal{R}}$ such that $Hu\not=Hv$;
  • is polynomial algebra, namely for any $R_{1},R_{2}\in\mathcal{H}_{\mathcal{R}}$ there exist $R^{\lambda}_{+},R_{\crossproduct}\in\mathcal{H}_{\mathcal{R}}$ such that $R_{1}(u)+\lambda R_{2}(u)=R_{+}^{\lambda}(u)\,and\quad R_{1}(u)\cdot R_{2}(u)=R_{\crossproduct}(u)$ for any $u\in\mathcal{I}$;
then $\mathcal{R}$ is universal for the set of mappings with fading memory.
Compactness of the input space
The mathematical proof of universality relies on the Stone-Weierestrass theorem, which applies to compact metric space. Thus, to ensure universality, we have to check that the input space, equipped with the fading metric introduced above is a compact metric space. This is a general result for time series inputs and it follows from the following theorem, proved also in [40].
Lemma IV.2.
The space of bounded time-dependent sequences $\mathcal{I}$ equipped with the distance $d_{\omega}$ defined in Def. I.1 is a compact metric space.
Echo state property and fading memory of the associated functional
A reservoir computer fulfills echo state property if the internal dynamics of the network does not depend explicitly on its specific initialization. This implies that the reservoir computer defines a unique, well-posed mapping between $\mathcal{I}$ and $\mathcal{O}$. As discussed above, both the echo state property and the fading memory follow if the dynamics is strictly contractive in the trace norm. Thus, these properties follow from the strict contractivity of the quantum channel that models the evolution of the echo state network, namely,
$\widetilde{\rho}_{t+1}=\mathcal{E}\left(U(u_{t+1})\rho_{t}U^{{\dagger}}(u_{t+1})\right)$
since the measurement process does not erase the strict contractivity of the overall evolution.
Separability of the inputs
A universal class of reservoir computers has the ability to discriminate any pair of different inputs. Namely, for any given $u,v\in\mathcal{I}$, the dynamics of a reservoir in the class should be such that associated mapping has different outputs when evaluated on $u$ and $v$, respectively. We proceed by constructing it, exploiting a single qubit reservoir and the action of suitable quantum channels. Denoting with $x\in[0,1]$ the current value of the input, the dynamics of reservoir register is written as
$\rho_{t+1}=\Phi_{x}(\rho_{t})=\mathcal{E}\left(U(x)\rho_{t}U^{\dagger}(x)\right)\,.$ (42)
Then, we can prove the following Theorem that ensures the separability of the input under suitable conditions on the dynamics.
Theorem IV.3.
If the dynamics described by Eq. (42) has a unique fixed point $\rho_{*}^{x}$, that depends univocally on $x$, then we have separability.
In fact, without loss of generality, let assume $u,v\in\mathcal{I}$ such that $u_{t}=v_{t}\,\,\forall t\not=0$ and $u_{0}=v_{0}$, and let denote with $\rho_{t}^{y},\,y=u,v$ the state of the circuit after the injection of $u$ and $v$, respectively. Since the input sequences are equal for $t<0$, we assume moreover that $\rho^{u}_{0}=\rho^{v}_{0}$. Preparing the circuit in the fixed point $\rho_{*}^{u_{0}}$, immediately we realize that $\rho_{1}^{u}\not=\rho_{1}^{v}$ since the fixed point $\rho_{1}^{u}=\rho_{*}^{u_{0}}$ depends univocally on $u_{0}$. Then the dynamics continues separated, as indeed the linear operator $e^{\mathcal{L}_{t}(y)}$, that encodes the evolution in Eq. (42), is a linear full rank operator.
It remains to show that there exists a quantum channel $\mathcal{E}$ such that Eq. (42) has a unique fixed point. This is the case for the amplitude damping channel.
Lemma IV.4.
If the quantum channel in Eq. (42) is the amplitude damping $\mathcal{E}_{\text{ad}}$, then the equation has a unique fixed point that depends univocally on the value of the input.
Proof.
It suffices to recall that the unique fixed point of the amplitude damping channel is the $|0\rangle$ state, namely $\mathcal{E}_{\text{ad}}(\rho)=\rho$ if and only if $\rho=|0\rangle\langle 0|$. Then, the fixed point of the composed map $\rho_{t+1}=\mathcal{E}_{\text{ad}}\left(U(x)\rho_{t}U^{\dagger}(x)\right)$ is simply $\rho_{*}^{x}=U^{\dagger}(x)|0\rangle\langle 0|U(x)$, which depends univocally on $x$. ∎
We remark that, on the other hand, if the quantum channel $\mathcal{E}$ is unital, namely if it preserves the identity density operator $\mathcal{E}\left(\mathbb{I}\right)=\mathbb{I}$, the dynamics has a fixed point that does not depend on the input values. Indeed, by definition of the unitality of a quantum channel, we have that
$\Phi_{x}(\rho_{*})=\rho_{*}\iff\rho_{*}=U(x)\mathbb{I}U^{\dagger}(x)=U(x)U^{\dagger}(x)\mathbb{I}=\mathbb{I}\,.$
regardless of the value of the input $x$.
Polynomial algebra structure by spatial multiplexing
It remains to prove that the associated mappings form a polynomial algebra, to ensure the universality of our family of quantum echo state networks. This can be achieved by exploiting spatial multiplexing, as widely known in the context of quantum reservoir computing [29, 18, 27]. It consists of parallelly preparing and running the two networks and taking the sum, or the product, of the readouts as the overall readout function. Then, for any two echo state networks whose associated mapping are respectively $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, it is possible to construct the systems whose associated mapping is the sum $\mathcal{C}_{1}+\mathcal{C}_{2}$ and the product $\mathcal{C}_{1}\cdot\mathcal{C}_{2}$. Adding this multiplexed architecture to the general class, we recover the structure of the polynomial algebra of the mappings associated. Thus, combining the previous Subsections, we have proved that, under suitable conditions on the quantum channel - for example, fulfilled by the amplitude damping channel - our class of gate-based echo state networks described is universal for the class of filters with fading memory.